17
Jan
2015
22:11

At approximately 19:52 this evening, one unit of a switch cluster failed in our core network. This resulted in outages of email, DNS and various other services. Functionality has been restored by remotely re-configuring core services to use other switch units, but connections provided directly by the failed switch remain down for now. We will announce planned maintenance to rectify this in due course.  We apologise for the inconvenience caused.

Posted by Kevin Hones

15
Jan
2015
14:49

Customers who have a TalkTalk connected service may have experienced a glitch in service for a minute or so at around 14:00 this afternoon, normal service is now running.

The cause of the glitch is being investigated by TalkTalk and as soon as we have more information we will post an update.

We apologise for any inconvenience caused.

Posted by Yolanda Zyczynska

10
Jan
2015
19:41

Update 22:02: We have received an explanation from our wholesaler as follows:
“At approximately 19:20 this evening there was a hardware failure on the talk talk dsl network. This affected dsl tails being delivered to us, the majority of which are now back online. However there are some sessions that are still down due to this failure. Talk Talk engineers are currently working to replace the failed card, but there is no eta for this work to be completed yet.. We will provide an update as soon as we get some more information.”

Unrelated to the earlier outage, a large number of TalkTalk lines went down at around 19:20. Connections are now returning. We suspect this is a problem within TalkTalk’s but have received no notifications yet from our wholesaler. We will post an update as and when we receive further information.

Posted by Kevin Hones

10
Jan
2015
09:59

Update 16:10: The link carrying TT services has now been restored. All services are now running again – if any routers don’t reconnect automatically, try power cycling and/or calling support.

Update 15:44: Services are finally coming back up – the link over which BT lines connect is now functioning. We are hopeful that TT lines will follow shortly.

Update 13:36: Our supplier has just stated “Please be advised that we are still experiencing an issue at Telehouse North. Engineers are currently onsite and actioning the solution.”

Update 11:10: The switch reboot apparently was unsuccessful. Our supplier has just advised that a replacement switch will be on site shortly.

There is currently a major problem with all DSL connections.
We have established that one of our suppliers currently has problems with a switch in Telehouse North and they are working on the problem.
When we hear more we will publish an update.

Posted by Kevin Hones

Current Status

    GoodNo current issues

Status Archive